Direct euro quote for pounds in Paris is 1.2345-468. Indirect euro quote for pounds in London is 0.7531-642. WHich currency to buy in which city for which price in which currency, which currency to sell in which city for which price in which currency and what is an arbitrage return?
Answer #1

Direct quote, domestic currency is price currency

Indirect quote, domestic currency is base currency

Pounds - GBP

Euro - EURO

GBP/EUR - GBP is price currency, EUR is base currency

EUR/GBP - EUR is price currency, GBP is base currency

So, GBP/EUR is indirect quote for EUR i.e. (direct quote for GBP) in Paris

(bid - ask) (1.2345-1.2468)

Also, EUR/GBP is indirect quote for GBP in London

(bid - ask) (0.7531-0.7642)

1 EUR can be bought in Paris for 1.2345 GBP. 1 GBP can be sold in Paris for 1/1.2345 i.e. 0.8100 EUR.

1 EUR can be sold in Paris for 1.2468 GBP. 1 GBP can be bought in Paris for 1/1.2468 i.e. 0.8021 EUR.

1 GBP can be bought in London for 0.7531 EUR.

1 GBP can be sold in London for 0.7642 EUR.

For Arbitrage, Buy LOW and Sell HIGH

It can be seen that, 1 GBP should be bought in London as it is undervalued as compared to Paris.

Step 1: Buy 1 GBP using 0.7531 EUR in London.

Step 2: Sell this 1 GBP in Paris for 0.81 EUR.

Arbitrage profit = 0.81 - 0.7531 = 0.0569 EUR

Arbitrage return = 0.0569/0.7531 = 7.56%

Please like. It took a lot of time and effort. Thanks.